OpenStudy (anonymous):

\[\int\limits_{0}^{x} xsin \Pi x\]

OpenStudy (anonymous):

there is something wrong here you cannot have the upper limit of integration by \(x\) and also the integrand by \(x\sin(\pi x)\)

OpenStudy (anonymous):

\[\int\limits_{0}^{x} xsin \Pi x\]

OpenStudy (anonymous):

but the second fundamental theorem of calculus applies

OpenStudy (anonymous):

\[\int_0^xx\sin(\pi x)dx\] makes no sense

OpenStudy (anonymous):

you could have \[\int_0^xt\sin(\pi t)dt\] for example

OpenStudy (anonymous):

or \[\int_0^{\pi}x\sin(\pi x)dx\] even

OpenStudy (anonymous):

okay maybe that's what the teacher meant

OpenStudy (anonymous):

can you post the exact question?

OpenStudy (anonymous):

that is it

OpenStudy (anonymous):

which one?

OpenStudy (anonymous):

\[\int\limits\limits_{0}^{x} xsin \Pi xdx\]